# FedPOD: the deployable units of training for federated learning

### v1

@@ Abstract
This paper proposes FedPOD (Proportionally Orchestrated Derivative) for optimizing learning efficiency and communication cost in federated learning among multiple clients. Inspired by FedPIDAvg [machler2022fedpidavg], we define a round-wise task for FedPOD to enhance training efficiency. FedPIDAvg achieved performance improvement by incorporating the training loss reduction for prediction entropy as weights using differential terms. Furthermore, by modeling data distribution with a Poisson distribution and using a PID controller, it reduced communication costs even in skewed data distribution. However, excluding participants classified as outliers based on the Poisson distribution can limit data utilization. Additionally, PID controller requires the same participants to be maintained throughout the federated learning process as it uses previous rounds’ learning information in the current round. In our approach, FedPOD addresses these issues by including participants excluded as outliers, eliminating dependency on previous rounds’ learning information, and applying a method for calculating validation loss at each round. In this challenge, FedPOD presents comparable performance to FedPIDAvg in metrics of Dice score, 0.78, 0.71 and 0.72 for WT, ET and TC in average, and projected convergence score, 0.74 in average. Furthermore, the concept of FedPOD draws inspiration from Kubernetes’ smallest computing unit, POD, designed to be compatible with Kubernetes auto-scaling. Extending round-wise tasks of FedPOD to POD units allows flexible design by applying scale-out similar to Kubernetes’ auto-scaling. This work demonstrated the potentials of FedPOD to enhance federated learning by improving efficiency, flexibility, and performance in metrics.
@@ 1.2 FeTS Challenge 2024
The FeTS challenge [pati2021federated, pati2021federated, karargyris2023federated] focuses on training and evaluating models for Glioblastoma (GBM) tumor segmentation using federated learning. Utilizing 1,473 BraTS dataset, which include institutional source information, the challenge performs semantic segmentation on three regions (NCR, ET, ED) using multi-modal MRI pairs (T1, T1-contrast, T2, T2-flair) [bakas10segmentation, bakas10segmentation, bakas2017advancing, baid2021rsna]. The challenge explores novel approaches to server-side node selection, hyper-parameter tuning, and aggregation methods to optimize federated learning. Model architecture and client-side optimization are uniformly applied to all participants, enabling an objective comparison of federated learning performance optimization.

### v2

@@ Abstract
This paper proposes FedPOD, which ranked first in the 2024 Federated Tumor Segmentation (FeTS) Challenge, for optimizing learning efficiency and communication cost in federated learning among multiple clients. Inspired by FedPIDAvg, we define a round-wise task for FedPOD to enhance training efficiency. FedPIDAvg achieved performance improvement by incorporating the training loss reduction for prediction entropy as weights using differential terms. Furthermore, by modeling data distribution with a Poisson distribution and using a PID controller, it reduced communication costs even in skewed data distribution. However, excluding participants classified as outliers based on the Poisson distribution can limit data utilization. Additionally, PID controller requires the same participants to be maintained throughout the federated learning process as it uses previous rounds’ learning information in the current round. In our approach, FedPOD addresses these issues by including participants excluded as outliers, eliminating dependency on previous rounds’ learning information, and applying a method for calculating validation loss at each round. In this challenge, FedPOD presents comparable performance to FedPIDAvg in metrics of Dice score, 0.78, 0.71 and 0.72 for WT, ET and TC in average, and projected convergence score, 0.74 in average. Furthermore, the concept of FedPOD draws inspiration from Kubernetes’ smallest computing unit, POD, designed to be compatible with Kubernetes auto-scaling. Extending round-wise tasks of FedPOD to POD units allows flexible design by applying scale-out similar to Kubernetes’ auto-scaling. This work demonstrated the potentials of FedPOD to enhance federated learning by improving efficiency, flexibility, and performance in metrics.
@@ 1.2 FeTS Challenge 2024
The Federated Tumor Segmentation (FeTS) challenge represents a premier international benchmark for federated learning in medical imaging, published in Nature Communications [zenk2025towards] and organized under the auspices of MICCAI, the leading conference in medical image computing [pati2021federated, karargyris2023federated]. This work presents FedPOD, our first-place solution [linardos2025miccai], which demonstrates state-of-the-art performance through efficient node selection and adaptive aggregation. The challenge focuses on training and evaluating models for Glioblastoma (GBM) tumor segmentation using federated learning across 1,473 BraTS dataset samples with institutional source information. The task involves semantic segmentation of three tumor regions (NCR, ET, ED) using multi-modal MRI pairs (T1, T1-contrast, T2, T2-flair) [bakas2017advancing, baid2021rsna]. The challenge explores novel approaches to server-side node selection, hyper-parameter tuning, and aggregation methods to optimize federated learning, while model architecture and client-side optimization are uniformly applied to all participants, enabling objective comparison of federated learning performance optimization.